OpenStudy (anonymous):

Find the reciprocal of the probability that a randomly chosen positive factor of 900 is less than 10. I need help on how to get started on this problem.

OpenStudy (anonymous1):

First, you need to find the positive factors of 900.

OpenStudy (anonymous1):

That is, find all the numbers that divide 900.

OpenStudy (anonymous1):

For example, 1 divides 900, because 900/1 = 900 (an integer). 2 also divides 900, because 900/2 = 450 (an integer). So, 1 and 2 are some of the positive factors of 900.